OpenStudy (anonymous):

\[35\sqrt{50}/7\sqrt{2}\]

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (anonymous):

the numerator can be simplified to 35sqrt50 = 35*sqrt(25*2) = 35*5*sqrt2 so (35*5*sqrt2)/(7*sqrt2) = 25 notice the sqrt2 cancels, and the 35 and 7 cancels 5/1 which leaves 5*5 at the numerator...

OpenStudy (anonymous):

will you check out my question plz? http://openstudy.com/study#/updates/4fc40f0de4b0964abc86ad6c

OpenStudy (anonymous):

\[(5\sqrt{2})(4\sqrt{10})\]